# Credal Prediction based on Relative Likelihood

### v1

@@ 1 Introduction
The distinction between two types of uncertainty, referred to as aleatoric and epistemic, is receiving increasing interest in machine learning (Hüllermeier and Waegeman, 2021). Roughly speaking, the aleatoric uncertainty of a predictive model is caused by the inherent randomness of the data-generating process, whereas epistemic uncertainty is caused by the learner’s lack of knowledge about the true (or best) predictive model. While aleatoric uncertainty is irreducible, epistemic uncertainty can be reduced on the basis of additional information, e.g., by collecting more training data.
@@
Aleatoric uncertainty can be captured adequately in terms of (conditional) probability distributions, whereas the representation of epistemic uncertainty requires “second-order” formalisms more general than (single) probability distributions. In the Bayesian approach to machine learning, second-order probability distributions are used for this purpose (Depeweg et al., 2018; Kendall and Gal, 2017): The learner maintains a probability distribution over the model space, which, if each model is a probabilistic predictor, results in predictions in the form of probability distributions of probability distributions (on outcomes). The Bayesian approach is theoretically appealing but computationally demanding, and commonly criticized for the need to specify a prior distribution (strongly influencing prediction and uncertainty quantification) (Gawlikowski et al., 2023).
@@
Figure 1: Overview of our method. Language inference example from the ChaosNLI dataset, showing a premise-hypothesis pair with three possible labels: Entailed, Neutral and Contradicted. Middle: Illustration of the learning mechanism of our Credal Relative Likelihood (CreRLα) framework, showing ensemble member training up to different $\alpha$ thresholds. Right: Unlike baselines, which are centered around the MLE and remain fixed, our method allows to adapt credal set size to improve coverage of the ground-truth distribution.
@@
An alternative second-order formalism is offered by credal sets (Walley, 1991), i.e., sets (instead of distributions) of probability distributions that are commonly assumed to be closed and convex (Cozman, 2000). Although a set appears to provide weaker information than a distribution, a set-based representation also has advantages. In particular, it can be argued that sets are more apt at representing ignorance in the sense of a lack of knowledge (Dubois et al., 1996), because distributions always involve additional assumptions beyond the mere distinction between plausible and implausible candidate models. Decision-making on the basis of credal sets has been proposed as a reasonable alternative to Bayesian decision-making (Levi, 1978; Girón and Ríos, 1980). The learning of credal predictors (i.e., models making predictions in the form of credal sets) has been considered in machine learning in the past (Zaffalon, 2001; Corani and Zaffalon, 2008) and, in light of the quest for proper representation of epistemic uncertainty, received renewed interest more recently (Wang et al., 2024a; Nguyen et al., 2025; Cella and Martin, 2024).
@@
A major question to be addressed in this regard concerns the distinction between plausible and implausible models. In this paper, we adopt a theoretically grounded approach for constructing credal sets based on the notion of relative likelihood. Our approach explicitly defines the prediction target as the set of all probability distributions generated by plausible learners, namely those whose relative likelihood exceeds a specified threshold (cf. Section 3). That said, realizing this in complex machine learning settings comes with several practical challenges, notably the questions of representation, approximation, and inference: How to formally represent subsets of a high-dimensional model space, how to approximate such sets algorithmically, and how to infer credal predictions from them? We propose methods to address these questions based on suitably modified ensemble techniques (cf. Section 4). We highlight the effectiveness of our approach on real-world datasets, evaluating its performance in terms of coverage and efficiency as key criteria of credal prediction. Also, we analyze its performance on downstream tasks such as Out-of-Distribution (OoD) detection (cf. Section 5).
Contributions. In summary our contributions are as follows:
@@
Related Work. In machine learning, uncertainty is often represented by (approximate) Bayesian methods (Mackay, 1992; Blundell et al., 2015; Lakshminarayanan et al., 2017; Gal and Ghahramani, 2016; Daxberger et al., 2021). An important characteristic of such representations, especially for uncertainty tasks, is diversity (D’Angelo and Fortuin, 2021; Wood et al., 2023), which can be enforced by regularization (de Mathelin et al., 2023), varying hyper-parameters (Wenzel et al., 2020), or increasing diversity in representations (Lopes et al., 2022). Alternatively, credal sets have been used in the fields of imprecise probability and machine learning to represent model uncertainty (Zaffalon, 2001; Corani and Zaffalon, 2008; Corani and Mignatti, 2015). Such sets can be generated based on the relative likelihood, also referred to as normalized likelihood (Antonucci et al., 2012a) and have been used in machine learning with simple model classes (Senge et al., 2014; Cella and Martin, 2024). In this work, we build upon these approaches and address the challenges that emerge when adapting the relative likelihood to a setting with complex predictors. Recently, credal sets have also been applied in the context of deep learning. Some approaches use ensemble learning to derive class-wise lower and upper probabilities (Wang et al., 2024b; Nguyen et al., 2025). Others train models to directly predict probability intervals (Wang et al., 2024a). Hybrid methods, combining multiple uncertainty frameworks, have also been proposed. Caprio et al. (2023) combine Bayesian deep learning and credal sets by considering sets of priors. Another approach leverages conformal prediction to construct credal sets with validity guarantees (Javanmardi et al., 2024). A more detailed discussion of the related work is provided in Appendix A.
@@ 2 Problem Statement
We consider classification in a supervised learning setting. We have access to training data $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is the instance space and $\mathcal{Y}=\{y_{1},\dots,y_{K}\}$ is the label space with $K\in\mathbb{N}$ classes. The training data are realizations of random variables that are independently and identically distributed according to a probability measure on $\mathcal{X}\times\mathcal{Y}$. We consider a hypothesis space $\mathcal{H}$ with probabilistic predictors $h:\mathcal{X}\to\Delta_{K}$, where $\Delta_{K}$ denotes the $(K-1)$-simplex (that we will also refer to as the probability simplex). Given an instance $\bm{x}$, a predictor $h$ assigns a probability distribution $p(\cdot\,|\,\bm{x},h)=h(\bm{x})$, which is an estimate of the ground-truth probability $p(\cdot\,|\,\bm{x},h^{*})$ generated by the ground-truth model $h^{*}$.
@@
We define the problem of learning a credal predictor as a generalization of the standard setting of supervised learning as introduced above: Given training data $\mathcal{D}$, the task is to induce a model $H:\mathcal{X}\rightarrow 2^{\Delta_{K}}$ that delivers predictions in the form of credal sets $\mathcal{Q}_{\bm{x}}=H(\bm{x})\subseteq\Delta_{K}$. In our setup, $H(\bm{x})$ represents the mapping of an instance $\bm{x}\in\mathcal{X}$ through a set of plausible models $\mathcal{C}$ to the corresponding credal set $\mathcal{Q}_{\bm{x}}$. In general, $H$ could also be realized differently. Inspired by other set-valued prediction methods such as conformal prediction (Vovk et al., 2005), we evaluate such a predictor in terms of its coverage and efficiency. Coverage (of the ground-truth distribution by the cedal set) is defined as follows:
@@
where $\overline{p}(y_{k}\,|\,\bm{x})=\sup\{p(y_{k}\,|\,\bm{x}):p\in H(\bm{x})\}$ (and $\underline{p}(y_{k}\,|\,\bm{x})$ is defined analogously). We assume efficiency to be positively oriented, meaning that higher efficiency corresponds to smaller credal sets. As opposed to the volume of a credal set, which, besides being challenging to compute, is not intuitive in high dimensions, this measure of efficiency is particularly interpretable. Specifically, it describes the (complement of the) average interval length for each class. Importantly, we use efficiency strictly as an indicator of set size, not as a measure of epistemic uncertainty (cf. Section 2).
@@ 3 Relative Likelihood-Based Credal Sets
Here, $h^{ML}$ represents the model in $\mathcal{H}$ with the highest likelihood given the training data $\mathcal{D}$. This notion, also called extended or normalized likelihood, has been proposed by Birnbaum (1962) and used for statistical inference (Wasserman, 1990; Walley and Moral, 1999). It offers an attractive alternative to Bayesian or frequentist reasoning eliminating the need to specify priors or perform resampling (Giang and Shenoy, 2002). In machine learning this is especially useful as specifying a meaningful prior on $\mathcal{H}$ is a non-trivial problem and resampling from the data distribution is usually not possible either, as we only have access to the given samples $\mathcal{D}$. An axiomatic justification of the relative likelihood (in the context of evidence theory) has been given by Denoeux (2014), who derives it from (i) the likelihood principle, (ii) compatibility with the Bayes rule, and (iii) the minimal commitment principle.
On the basis of the relative likelihood, a set of models can be constructed, analogous to a confidence region (Aitkin, 1982), by including models that are plausible in the sense of surpassing a threshold $\alpha\in[0,1]$. This set of models is also referred to as an $\alpha$-cut (Antonucci et al., 2012a):
@@
Based on 10 samples from a Bernoulli distribution with parameter 0.5. The credal set of $\alpha=0.9$ around the MLE at $0.6$ does not cover the ground-truth. Lowering $\alpha$ increases coverage at the cost of efficiency.
@@
According to this definition, a model $h$ is considered implausible (and hence ignored) if its likelihood is too small compared to the likelihood of the (presumably) best model, namely if its likelihood is less than $\alpha$ times the likelihood of the best model. Thus, the parameter $\alpha$ has a quite intuitive meaning. The concept of $\alpha$-cuts is further illustrated in Figure 2.
@@
Considering that we aim to simultaneously optimize coverage and efficiency, it is important to highlight the role of the parameter $\alpha$ as a means to move along the Pareto front and trade coverage against efficiency: The lower $\alpha$, the higher coverage tends to be. At the same time, however, efficiency will deteriorate (cf. again Figure 2). As $\alpha$ increases, the intervals for given data points shrink since only models with high relative likelihood for the given data sample are included. Conversely, lowering $\alpha$ allows for the inclusion of models with lower likelihood, resulting in larger credal intervals. This controllability induced by $\alpha$ emphasizes the appeal of the relative likelihood as a tool for (plausible) model selection and, thereby, credal set construction.
@@ 4 Approximating Relative Likelihood-Based Credal Sets
Estimating the Maximum Likelihood Estimator. Several strategies can be employed to obtain $h^{ML}$, including training a model with parameters known to perform well, utilizing a pre-trained model, or leveraging AutoML techniques to determine optimal parameter settings during training. Throughout this work, we adopt the first approach and train the maximum likelihood estimator $h^{ML}$ using parameter configurations that have been identified in the literature as effective. Further details and configurations of experiments are provided in Appendix B.
@@
Approximating the $\bm{\alpha}$-cut. Having found a suitable candidate model to compute the maximum likelihood, the next step is to construct the $\alpha$-cut. A straightforward approach is to train an ensemble of hypotheses and construct the $\alpha$-cut based on these hypotheses. In essence, approaches based on regular ensemble training (Wang et al., 2024b; Nguyen et al., 2025) form an example of this strategy. However, an issue is that all hypotheses $h$ typically tend to cluster around $h^{ML}$, hence the obtained hypotheses do not accurately approximate the $\alpha$-cut (unless $\alpha\approx 1$, see Figure 1).
To obtain better coverage of $\mathcal{C}_{\alpha}$, we propose to train $M$ hypotheses using an early stopping strategy, namely until a specific relative likelihood value $\tau_{i}$ is reached. Given $\alpha$, we consider thresholds $\tau=\{\tau_{i}\,|\,\tau_{i}=\alpha+i\cdot\Delta\tau\}$, where $\Delta\tau=\frac{1-\alpha}{M-1}$ and $i\in\{0,\ldots,M-2\}$. This guarantees a broad range of hypotheses in terms of relative likelihoods. The influence of the ensemble size $M$ is analyzed in an ablation study in Section C.2.
Although this should cover the $\alpha$-cut well, there might still be a lack of diversity in the predictions of the resulting hypotheses. In the literature, it is known that training ensembles without explicitly encouraging diversity actually minimizes diversity (Abe et al., 2022). To encourage diversity, we introduce a novel initialization strategy called ToBias. The idea is to make sure that the initial state of the ensemble, i.e., prior to training, represents a state of full uncertainty (or no knowledge). Hence, the resulting credal prediction should entail the entire probability simplex. As the predictors in the ensemble are trained, the amount of knowledge increases, and the predicted credal set should shrink. To enforce this in a finite predictor scenario, we ensure that the initial predictions of the learners correspond to degenerate probability distributions at vertices of the $(K-1)$-simplex, effectively covering the entire simplex in the initially predicted credal set. Figure 1 illustrates the mechanism of ToBias initialization and the overall learning process.
Technically, the initialization is performed by assigning a large constant to one of the biases in the final layer of each predictor. For example, given the biases $\bm{b}_{i}=[b_{i,1},...,b_{i,K}]$ of the last layer of predictor $h_{i}$, we set $b_{i,\ i\ \text{mod}\ K}=\beta$, where $\beta$ is large constant, following the regular initialization. In the remainder of this work, we use $\beta=100$. An ablation study examining the influence of ToBias initialization is presented in Section C.2.
@@
We illustrate the difference between regular initialization and ToBias initialization in Figure 3. Observe that without ToBias initialization, the credal set initially represents a state of low epistemic uncertainty, as all predictions are concentrated around the barycenter of the probability simplex, and increases its uncertainty as training progresses. The opposite can be observed with ToBias initialization. The credal set starts representing a state of full uncertainty and decreases its uncertainty as the learners acquire more knowledge.
@@
where $\hat{\gamma}(h_{i})\approx\tau_{i}$ is enforced by the training process described above. The full algorithm to train Credal Relative Likelihood (CreRLα) ensembles is presented in Algorithm 1.
Credal Set Predictions. Given a query instance $\bm{x}\in\mathcal{X}$, the set $\tilde{\mathcal{C}}_{\alpha}$ induces a finite approximation of the credal set:
@@
There are several principled ways of transforming this finite set into a credal set. Two common approaches are to derive (i) the convex hull of the set $\tilde{\mathcal{Q}}_{\bm{x},\alpha}$ and (ii) the intervals (lower and upper probabilities) for individual classes obtained from this set. Note that the convex hull is contained in the “box” credal set induced by the intervals. In the remainder of this work, we opt for the latter approach and define the approximation of $\mathcal{Q}_{\bm{x},\alpha}$ as follows (de Campos et al., 1994):

### v2

@@ 1 Introduction
The distinction between two types of uncertainty, referred to as aleatoric and epistemic, is receiving increasing interest in machine learning (hullermeier2021aleatoric). Roughly speaking, the aleatoric uncertainty of a predictive model is caused by the inherent randomness of the data-generating process, whereas epistemic uncertainty is caused by the learner’s lack of knowledge about the true (or best) predictive model. While aleatoric uncertainty is irreducible, epistemic uncertainty can be reduced on the basis of additional information, e.g., by collecting more training data. ***equal contribution
@@
Aleatoric uncertainty can be captured adequately in terms of (conditional) probability distributions, whereas the representation of epistemic uncertainty requires “second-order” formalisms more general than (single) probability distributions. In the Bayesian approach to machine learning, second-order probability distributions are used for this purpose (depewegDecompositionUncertainty2018; kendall2017uncertainties): The learner maintains a probability distribution over the model space, which, if each model is a probabilistic predictor, results in predictions in the form of probability distributions of probability distributions (on outcomes). The Bayesian approach is theoretically appealing but computationally demanding, and commonly criticized for the need to specify a prior distribution (strongly influencing prediction and uncertainty quantification) (gawlikowskiSurveyUncertainty2023).
@@
Figure 1: Overview of our method. Language inference example from the ChaosNLI dataset, showing a premise-hypothesis pair with three possible labels: Entailed, Neutral and Contradicted. Middle: Illustration of the learning mechanism of our Credal Relative Likelihood (CreRL) framework, showing ensemble member training up to different $\alpha$ thresholds. Right: Unlike baselines, which are centered around the MLE and remain fixed, our method allows to adapt credal set size to improve coverage of the ground-truth distribution.
@@
An alternative second-order formalism is offered by credal sets (walleyStatisticalReasoning1991), i.e., sets (instead of distributions) of probability distributions that are commonly assumed to be closed and convex (cozman2000credal). Although a set appears to provide weaker information than a distribution, a set-based representation also has advantages. In particular, it can be argued that sets are more apt at representing ignorance in the sense of a lack of knowledge (dubo_rp96), because distributions always involve additional assumptions beyond the mere distinction between plausible and implausible candidate models. Decision-making on the basis of credal sets has been proposed as a reasonable alternative to Bayesian decision-making (leviIndeterminateProbabilities1978; gironQuasiBayesian1980). The learning of credal predictors (i.e., models making predictions in the form of credal sets) has been considered in machine learning in the past (zaffalonStatisticalInference2001; coraniLearningReliable2008) and, in light of the quest for proper representation of epistemic uncertainty, received renewed interest more recently (wangCredalDeep2024; nguyenCredalEnsembling2025; cellaVariationalApproximations2024).
@@
A major question to be addressed in this regard concerns the distinction between plausible and implausible models. In this paper, we adopt a theoretically grounded approach for constructing credal sets based on the notion of relative likelihood. Our approach explicitly defines the prediction target as the set of all probability distributions generated by plausible learners, namely those whose relative likelihood exceeds a specified threshold (cf. Section˜3). That said, realizing this in complex machine learning settings comes with several practical challenges, notably the questions of representation, approximation, and inference: How to formally represent subsets of a high-dimensional model space, how to approximate such sets algorithmically, and how to infer credal predictions from them? We propose methods to address these questions based on suitably modified ensemble techniques (cf. Section˜4). We highlight the effectiveness of our approach on real-world datasets, evaluating its performance in terms of coverage and efficiency as key criteria of credal prediction. Also, we analyze its performance on downstream tasks such as Out-of-Distribution (OoD) detection (cf. Section˜5).
Contributions.
In summary our contributions are as follows:
@@
Related Work.
In machine learning, uncertainty is often represented by (approximate) Bayesian methods (mackayBayesianMethods1992; blundellWeightUncertainty2015; lakshminarayananDeepEnsembles2017; galDropoutAs2016; daxbergerLaplaceRedux2021). An important characteristic of such representations, especially for uncertainty tasks, is diversity (d2021repulsive; woodUnifiedTheory2023), which can be enforced by regularization (deMathelinDeepAnti2023), varying hyper-parameters (wenzelHyperparameterEnsembles2020), or increasing diversity in representations (lopesNoOne2022). Alternatively, credal sets have been used in the fields of imprecise probability and machine learning to represent model uncertainty (zaffalonStatisticalInference2001; coraniLearningReliable2008; coraniCredalModel2015). Such sets can be generated based on the relative likelihood, also referred to as normalized likelihood (antonucciLikelihoodBased2012) and have been used in machine learning with simple model classes (sengeReliableClassification2014; cellaVariationalApproximations2024). In this work, we build upon these approaches and address the challenges that emerge when adapting the relative likelihood to a setting with complex predictors. Recently, credal sets have also been applied in the context of deep learning. Some approaches use ensemble learning to derive class-wise lower and upper probabilities (wangCredalWrapper2024; nguyenCredalEnsembling2025). Others train models to directly predict probability intervals (wangCredalDeep2024). Hybrid methods, combining multiple uncertainty frameworks, have also been proposed. caprioImpreciseBayesian2023 combine Bayesian deep learning and credal sets by considering sets of priors. Another approach leverages conformal prediction to construct credal sets with validity guarantees (javanmardiConformalizedCredal2024). A more detailed discussion of the related work is provided in Appendix˜A.
@@ 2 Problem Statement
We consider classification in a supervised learning setting. We have access to training data $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is the instance space and $\mathcal{Y}=\{y_{1},\dots,y_{K}\}$ is the label space with $K\in\mathbb{N}$ classes. The training data are realizations of random variables that are independently and identically distributed according to a probability measure on $\mathcal{X}\times\mathcal{Y}$. We consider a hypothesis space $\mathcal{H}$ with probabilistic predictors $h:\mathcal{X}\to\Delta_{K}$, where $\Delta_{K}$ denotes the (K1)(K-1)-simplex (that we will also refer to as the probability simplex). Given an instance $\bm{x}$, a predictor $h$ assigns a probability distribution $p(\cdot\,|\,\bm{x},h)=h(\bm{x})$, which is an estimate of the ground-truth probability $p(\cdot\,|\,\bm{x},h^{*})$ generated by the ground-truth model $h^{*}$.
@@
We define the problem of learning a credal predictor as a generalization of the standard setting of supervised learning as introduced above: Given training data $\mathcal{D}$, the task is to induce a model $H:\mathcal{X}\rightarrow 2^{\Delta_{K}}$ that delivers predictions in the form of credal sets $\mathcal{Q}_{\bm{x}}=H(\bm{x})\subseteq\Delta_{K}$. In our setup, $H(\bm{x})$ represents the mapping of an instance $\bm{x}\in\mathcal{X}$ through a set of plausible models $\mathcal{C}$ to the corresponding credal set $\mathcal{Q}_{\bm{x}}$. In general, $H$ could also be realized differently. Inspired by other set-valued prediction methods such as conformal prediction (vovk2005algorithmic), we evaluate such a predictor in terms of its coverage and efficiency. Coverage (of the ground-truth distribution by the cedal set) is defined as follows:
@@
where $\overline{p}(y_{k}\,|\,\bm{x})=\sup\{p(y_{k}\,|\,\bm{x}):p\in H(\bm{x})\}$ (and $\underline{p}(y_{k}\,|\,\bm{x})$ is defined analogously). We assume efficiency to be positively oriented, meaning that higher efficiency corresponds to smaller credal sets. As opposed to the volume of a credal set, which, besides being challenging to compute, is not intuitive in high dimensions, this measure of efficiency is particularly interpretable. Specifically, it describes the (complement of the) average interval length for each class. Importantly, we use efficiency strictly as an indicator of set size, not as a measure of epistemic uncertainty (cf. Section˜2).
@@ 3 Relative Likelihood-Based Credal Sets
Here, $h^{ML}$ represents the model in $\mathcal{H}$ with the highest likelihood given the training data $\mathcal{D}$. This notion, also called extended or normalized likelihood, has been proposed by birnbaumFoundationsStatistical1962 and used for statistical inference (wassermanBeliefFunctions1990; walleyUpperProbabilities1999). It offers an attractive alternative to Bayesian or frequentist reasoning eliminating the need to specify priors or perform resampling (giangStatisticalDecisions2002). In machine learning this is especially useful as specifying a meaningful prior on $\mathcal{H}$ is a non-trivial problem and resampling from the data distribution is usually not possible either, as we only have access to the given samples $\mathcal{D}$. An axiomatic justification of the relative likelihood (in the context of evidence theory) has been given by denoeuxLikelihoodBased2014, who derives it from (i) the likelihood principle, (ii) compatibility with the Bayes rule, and (iii) the minimal commitment principle.
@@
Based on 10 samples from a Bernoulli distribution with parameter 0.5. The credal set of $\alpha=0.9$ around the MLE at $0.6$ does not cover the ground-truth. Lowering $\alpha$ increases the coverage at the cost of efficiency.
On the basis of the relative likelihood, a set of models can be constructed, analogous to a confidence region (aitkinDirectLikelihood1982), by including models that are plausible in the sense of surpassing a threshold $\alpha\in[0,1]$. This set of models is also referred to as an $\alpha$-cut (antonucciLikelihoodBased2012):
@@
According to this definition, a model $h$ is considered implausible (and hence ignored) if its likelihood is too small compared to the likelihood of the (presumably) best model, namely if its likelihood is less than $\alpha$ times the likelihood of the best model. Thus, the parameter $\alpha$ has a quite intuitive meaning. The concept of $\alpha$-cuts is illustrated in Figure˜2.
@@
Considering that we aim to simultaneously optimize coverage and efficiency, it is important to highlight the role of the parameter $\alpha$ as a means to move along the Pareto front and trade coverage against efficiency: The lower $\alpha$, the higher coverage tends to be. At the same time, however, efficiency will deteriorate (cf. again Figure˜2). As $\alpha$ increases, the intervals for given data points shrink since only models with high relative likelihood for the given data sample are included. Conversely, lowering $\alpha$ allows for the inclusion of models with lower likelihood, resulting in larger credal intervals. This controllability induced by $\alpha$ emphasizes the appeal of the relative likelihood as a tool for (plausible) model selection and, thereby, credal set construction.
@@ 4 Approximating Relative Likelihood-Based Credal Sets
Estimating the Maximum Likelihood Estimator.
Several strategies can be employed to obtain $h^{ML}$, including training a model with parameters known to perform well, utilizing a pre-trained model, or leveraging AutoML techniques to determine optimal parameter settings during training. Throughout this work, we adopt the first approach and train the maximum likelihood estimator $h^{ML}$ using parameter configurations that have been identified in the literature as effective. Further details and configurations of experiments are provided in Appendix˜B.
@@
Approximating the $\bm{\alpha}$-cut.
Having found a suitable candidate model to compute the maximum likelihood, the next step is to construct the $\alpha$-cut. A straightforward approach is to train an ensemble of hypotheses and construct the $\alpha$-cut based on these hypotheses. In essence, approaches based on regular ensemble training (wangCredalWrapper2024; nguyenCredalEnsembling2025) form an example of this strategy. However, an issue is that all hypotheses $h$ typically tend to cluster around $h^{ML}$, hence the obtained hypotheses do not accurately approximate the $\alpha$-cut (unless $\alpha\approx 1$, see Figure˜1).
To obtain better coverage of $\mathcal{C}_{\alpha}$, we propose to train $M$ hypotheses using an early stopping strategy, namely until a specific relative likelihood value $\tau_{i}$ is reached. Given $\alpha$, we consider thresholds $\tau=\{\tau_{i}\,|\,\tau_{i}=\alpha+i\cdot\Delta\tau\}$, where $\Delta\tau=\frac{1-\alpha}{M-1}$ and $i\in\{0,\ldots,M-2\}$. This guarantees a broad range of hypotheses in terms of relative likelihoods. The influence of the ensemble size $M$ is analyzed in an ablation study in Section˜C.2.
Although this should cover the $\alpha$-cut well, there might still be a lack of diversity in the predictions of the resulting hypotheses. In the literature, it is known that training ensembles without explicitly encouraging diversity actually minimizes diversity (abeBestDeep2022). To encourage diversity, we introduce a novel initialization strategy called ToBias. The idea is to make sure that the initial state of the ensemble, i.e., prior to training, represents a state of full uncertainty (or no knowledge). Hence, the resulting credal prediction should entail the entire probability simplex. As the predictors in the ensemble are trained, the amount of knowledge increases, and the predicted credal set should shrink. To enforce this in a finite predictor scenario, we ensure that the initial predictions of the learners correspond to degenerate probability distributions at vertices of the (K1)(K-1)-simplex, effectively covering the entire simplex in the initially predicted credal set. Figure˜1 illustrates the mechanism of ToBias initialization and the overall learning process.
Technically, the initialization is performed by assigning a large constant to one of the biases in the final layer of each predictor. For example, given the biases $\bm{b}_{i}=[b_{i,1},...,b_{i,K}]$ of the last layer of predictor $h_{i}$, we set $b_{i,\ i\ \text{mod}\ K}=\beta$, where $\beta$ is large constant, following the regular initialization. In the remainder of this work, we use $\beta=100$. An ablation study examining the influence of ToBias initialization is presented in Section˜C.2.
@@
We illustrate the difference between regular initialization and ToBias initialization in Figure˜3. Observe that without ToBias initialization, the credal set initially represents a state of low epistemic uncertainty, as all predictions are concentrated around the barycenter of the probability simplex, and increases its uncertainty as training progresses. The opposite can be observed with ToBias initialization. The credal set starts representing a state of full uncertainty and decreases its uncertainty as the learners acquire more knowledge.
@@
where $\hat{\gamma}(h_{i})\approx\tau_{i}$ is enforced by the training process described above. The full algorithm to train Credal Relative Likelihood (CreRL) ensembles is presented in Algorithm˜1.
Credal Set Predictions.
Given a query instance $\bm{x}\in\mathcal{X}$, the set $\tilde{\mathcal{C}}_{\alpha}$ induces a finite approximation of the credal set:
@@
There are several principled ways of transforming this finite set into a credal set. Two common approaches are to derive (i) the convex hull of the set $\tilde{\mathcal{Q}}_{\bm{x},\alpha}$ and (ii) the intervals (lower and upper probabilities) for individual classes obtained from this set. Note that the convex hull is contained in the “box” credal set induced by the intervals. In the remainder of this work, we opt for the latter approach and define the approximation of $\mathcal{Q}_{\bm{x},\alpha}$ as follows (camposProbabilityIntervals1994):